# LongReD: Mitigating Short-Text Degradation of Long-Context Large Language Models via Restoration Distillation

### v2

@@ 1 Introduction
Large language models (LLMs) have exhibited remarkable performance across a wide range of text and multimodal tasks Brown et al. (2020); OpenAI (2023); Zhao et al. (2023); Touvron et al. (2023); Dubey et al. (2024). However, their abilities to process long contexts are constrained by the positional encodings and the attention mechanism, which define the context window size based on the length of the pre-training data Su et al. (2024); Press et al. (2022). When the input text exceeds this context window, the model encounters out-of-distribution positional information, resulting in significant performance degradation Chen et al. (2023); Peng et al. (2023b); Dong et al. (2024).
To address the challenges of long context processing, various methods have been proposed to extend the context window of LLMs bloc97 (2023); Xiao et al. (2023); Xiong et al. (2024); Chen et al. (2023). A prominent approach involves scaling positional encodings combined with lightweight continual pre-training. Leveraging the properties of RoPE Su et al. (2024), these methods interpolate positional indices or adjust the RoPE base to prevent out-of-distribution rotary angles beyond the original context window Chen et al. (2023); Ding et al. (2024); Peng et al. (2023b); Xiong et al. (2024). Subsequently, lightweight continual pre-training adapts LLMs to the extended context and modified encodings, enabling context window up to 128K or even 1M tokens Xiong et al. (2024); Fu et al. (2024); Zeng et al. (2024). However, as shown in Figure 1, these long context window extension techniques come at the cost of degraded performance on short-text tasks Xiong et al. (2024). Yet, the root causes of this performance decline and potential mitigation strategies remain under-explored.
@@ 2.1 Transformer
Owing to strong capabilities and scalability, Transformer decoders have served as the backbone of most LLMs Vaswani et al. (2017); Touvron et al. (2023); Dubey et al. (2024). Given a Transformer decoder with $L$ layers and $N$ heads and an input sequence $\bm{x}=\{x_{1},\dots,x_{T}\}$ consisting with $T$ tokens, the output of the $l$-th layer can be denoted as $\mathbf{H}_{l}=\{\bm{h}_{l,1},\dots,\bm{h}_{l,T}\}$. Each Transformer layer consists of a multi-head attention (MHA) module and a feed-forward network (FFN) module with residual connections connecting them, as shown by the following formula:
@@ 4.2 Short-Text Distillation
Distillation Layer Selection. To decrease the distribution discrepancy, we choose to distill the inner hidden states of LLMs. Yet, ensuring all the hidden states are the same as the original model is impossible and may significantly decrease the long text modeling capacities. Thus, we only select $M(<L)$ layers as the distill layers $\mathbf{L}_{M}=\{l_{1},\dots,l_{M}\},\forall l_{i}\in\{0,\dots,L\}$ and distill the outputs of these layers. Specifically, we propose an attention-based layer selection method where we first compute the KL divergence of attention scores between the extended and original models and then select layers with the largest KL divergences for distillation. We also select the critical last layer that directly determines the output Men et al. (2024).
@@ 5.3 Detailed Analysis
Ablation Study. In addition to long-text training, our method incorporates both short-text distillation and short-to-long distillation. To evaluate the individual contributions of these two components, we employ half of the training data for long-text training and another half data for short-text distillation or short-to-long distillation. Moreover, we also evaluate the effect of the hyper-parameters $\alpha_{1}$ and $\alpha_{2}$ on both short and long text tasks. The results are compared against the LongReD-C method with $\alpha_{1}=5,\alpha_{2}=10$, as summarized in Table 4. The results indicate that excluding short-text distillation significantly degrades the model’s performance on short-text tasks. Conversely, omitting the short-to-long distillation stage results in noticeable declines in long-text performance. In addition, The adaptation of the hyper-parameters is essential. Increasing or decreasing $\alpha_{2}$ will lead to performance degradation on long-text tasks, while slightly decreasing $\alpha_{1}$ will lead to better long-text performances. When either $\alpha_{1}$ or $\alpha_{2}$ is set to an extremely large value, the model’s performance on long-context tasks drops precipitously.
@@
Comparison with Continual Learning Methods. Beyond simple continual pre-training, we also compare LongReD with two continual learning methods that can effectively mitigate catastrophic forgetting issues: (1) Model Merging: Following previous work Hu et al. (2024b), we average the checkpoints of the original model and the extended model; (2) Parameter-Efficient Tuning: Since modifying RoPE directly affects attention computation without altering the knowledge stored in MLP parameters, we only tune the parameters in the attention blocks. We train Llama-3-8B on a 32K context length with these methods and evaluate their performances. As shown in Table 7, though these methods can achieve better performances than naive continual pre-training, our method consistently performs better on most short-context capacities evaluations.
@@ 6 Related Work
Context Window Extension. LLMs typically have constrained context windows owing to positional encodings and attention mechanisms Press et al. (2022); Han et al. (2023). In order to meet the growing demand for processing long texts, various methods have been proposed to extend the context windows of LLMs based on modification of RoPE Su et al. (2024). Positional Interpolation down-scale positional indices to prevent out-of-distribution positions. In addition, NTK bloc97 (2023); emozilla (2023), ABF Xiong et al. (2024), Yarn Peng et al. (2023b), and LongRoPE Ding et al. (2024) modify the base of RoPE to control the rotation angles of different dimensions. After modifying positional encodings, LLMs are lightweightly continued pre-training on long texts Fu et al. (2024). Through this pipeline, the context windows of LLMs can be extended to 128K or even 1M tokens Dubey et al. (2024); Zeng et al. (2024). Typically, they can achieve utilization of long texts within their context windows at the cost of degradation of general capacities on short texts Dubey et al. (2024). Our work mainly focuses on the reasons for the degradation of short-text capacities and approaches to mitigating the decline.
@@ Appendix E Analysis of Distillation Length
As discussed in Section 5.3, increasing the distillation length will harm the long text modeling capacities. To explore the reason for the performance decline, we employ positional vectors Dong et al. (2024), a method to extract latent positional information from the hidden states. Specifically, given the hidden states from different samples, we average the hidden states at the same position to obtain vectors that representing positional information, denoted as positional vectors:
$\bm{p}_{l,i}=\frac{1}{N}\sum_{s=1}^{N}\bm{h}_{l,i}^{s}$ (40)
Appendix F Results Details

### v3

@@ 1 Introduction
Large language models (LLMs) have exhibited remarkable performance across a wide range of text and multimodal tasks Brown et al. (2020); OpenAI (2023); Zhao et al. (2023); Touvron et al. (2023); Dubey et al. (2024); Tang et al. (2025c); Wang et al. (2025); Chen et al. (2025). However, their abilities to process long contexts are constrained by the positional encodings and the attention mechanism, which define the context window size based on the length of the pre-training data Su et al. (2024); Press et al. (2022). When the input text exceeds this context window, the model encounters out-of-distribution positional information, resulting in significant performance degradation Chen et al. (2023); Peng et al. (2023b); Dong et al. (2024a).
To address the challenges of long context processing, various methods have been proposed to extend the context window of LLMs bloc97 (2023); Xiao et al. (2023); Xiong et al. (2024); Chen et al. (2023). A prominent approach involves scaling positional encodings combined with lightweight continual pre-training. Leveraging the properties of RoPE Su et al. (2024), these methods interpolate positional indices or adjust the RoPE base to prevent out-of-distribution rotary angles beyond the original context window Chen et al. (2023); Ding et al. (2024); Peng et al. (2023b); Xiong et al. (2024). Subsequently, lightweight continual pre-training adapts LLMs to the extended context and modified encodings, enabling context window up to 128K or even 1M tokens Xiong et al. (2024); Fu et al. (2024); Zeng et al. (2024). However, as shown in Figure 1, these long context window extension techniques come at the cost of degraded performance on short-text tasks Xiong et al. (2024); Dong et al. (2024b). Yet, the root causes of this performance decline and potential mitigation strategies remain under-explored.
@@ 2.1 Transformer
Owing to strong capabilities and scalability, Transformer decoders have served as the backbone of most LLMs Vaswani et al. (2017); Touvron et al. (2023); Dubey et al. (2024); Tang et al. (2025b). Given a Transformer decoder with $L$ layers and $N$ heads and an input sequence $\bm{x}=\{x_{1},\dots,x_{T}\}$ consisting with $T$ tokens, the output of the $l$-th layer can be denoted as $\mathbf{H}_{l}=\{\bm{h}_{l,1},\dots,\bm{h}_{l,T}\}$. Each Transformer layer consists of a multi-head attention (MHA) module and a feed-forward network (FFN) module with residual connections connecting them, as shown by the following formula:
@@ 4.2 Short-Text Distillation
Distillation Layer Selection. To decrease the distribution discrepancy, we choose to distill the inner hidden states of LLMs. Yet, ensuring all the hidden states are the same as the original model is impossible and may significantly decrease the long text modeling capacities. Thus, we only select $M(<L)$ layers as the distill layers $\mathbf{L}_{M}=\{l_{1},\dots,l_{M}\},\forall l_{i}\in\{0,\dots,L\}$ and distill the outputs of these layers. Specifically, we propose an attention-based layer selection method where we first compute the KL divergence of attention scores between the extended and original models and then select layers with the largest KL divergences for distillation Men et al. (2024); Dong et al. (2025). We also select the critical last layer that directly determines the output.
@@ 5.3 Detailed Analysis
Ablation Study. In addition to long-text training, our method incorporates both short-text distillation and short-to-long distillation. To evaluate the individual contributions of these two components, we employ half of the training data for long-text training and the other half for short-text distillation or short-to-long distillation. Moreover, we also evaluate the effect of the hyper-parameters $\alpha_{1}$ and $\alpha_{2}$ on both short and long text tasks. The results are compared against the LongReD-C method with $\alpha_{1}=5,\alpha_{2}=10$, as summarized in Table 4. The results indicate that excluding short-text distillation significantly degrades the model’s performance on short-text tasks. Conversely, omitting the short-to-long distillation stage results in noticeable declines in long-text performance. In addition, the adaptation of the hyper-parameters is essential. Increasing or decreasing $\alpha_{2}$ will lead to performance degradation on long-text tasks, while slightly decreasing $\alpha_{1}$ will lead to better long-text performances. When either $\alpha_{1}$ or $\alpha_{2}$ is set to an extremely large value, the model’s performance on long-context tasks drops precipitously.
@@
Comparison with Continual Learning Methods. Beyond simple continual pre-training, we also compare LongReD with two continual learning methods that can effectively mitigate catastrophic forgetting issues: (1) Model Merging: Following common model merging methods Hu et al. (2024b); Chen et al. (2024b), we average the checkpoints of the original model and the extended model; (2) Parameter-Efficient Tuning: Since modifying RoPE directly affects attention computation without altering the knowledge stored in MLP parameters, we only tune the parameters in the attention blocks Chen et al. (2024a). We train Llama-3-8B on a 32K context length with these methods and evaluate their performance. As shown in Table 7, though these methods can achieve better performances than naive continual pre-training, our method consistently performs better on most short-context capacities evaluations. This may be because these methods only address catastrophic forgetting. However, our approach not only addresses catastrophic forgetting through continual learning but also leverages knowledge distillation to alleviate distribution drift.
@@ 6 Related Work
Context Window Extension. LLMs typically have constrained context windows owing to positional encodings and attention mechanisms Press et al. (2022); Han et al. (2023); Dong et al. (2023). In order to meet the growing demand for processing long texts Wang et al. (2024); Tang et al. (2025a); Peng et al. (2025), various methods have been proposed to extend the context windows of LLMs based on modification of RoPE Su et al. (2024). Positional Interpolation down-scales positional indices to prevent out-of-distribution positions. In addition, NTK bloc97 (2023); emozilla (2023), ABF Xiong et al. (2024), Yarn Peng et al. (2023b), and LongRoPE Ding et al. (2024) modify the base of RoPE to control the rotation angles of different dimensions. After modifying positional encodings, LLMs are lightweightly continued pre-training on long texts Fu et al. (2024). Through this pipeline, the context windows of LLMs can be extended to 128K or even 1M tokens Dubey et al. (2024); Zeng et al. (2024). Typically, they can achieve utilization of long texts within their context windows at the cost of degradation of general capacities on short texts Dubey et al. (2024). Our work mainly focuses on the reasons for the degradation of short-text capacities and approaches to mitigating the decline.
@@ 7 Conclusion
Acknowledgement
This work was partially supported by National Natural Science Foundation of China under Grant No. 92470205 and 62222215, Beijing Municipal Science and Technology Project under Grant No. Z231100010323009, and Beijing Natural Science Foundation under Grant No. L233008. Xin Zhao is the corresponding author.
@@ Appendix C Evaluation Details
We evaluate these models using the LLMBox Tang et al. (2024) library for most benchmarks, and OpenCompass Contributors (2023) for the math-related benchmarks.
@@ Appendix E Analysis of Distillation Length
As discussed in Section 5.3, increasing the distillation length will harm the long text modeling capacities. To explore the reason for the performance decline, we employ positional vectors Dong et al. (2024a), a method to extract latent positional information from the hidden states. Specifically, given the hidden states from different samples, we average the hidden states at the same position to obtain vectors that representing positional information, denoted as positional vectors:
$\bm{p}_{l,i}=\frac{1}{N}\sum_{s=1}^{N}\bm{h}_{l,i}^{s}$ (40)
Appendix F Results Details